Write the point-slope form of the equation of the line described.

Through: (4,2), parallel to y=-3/4x-5

Answer by Alan3354(69443)

Step 1, find the slope of the given line.
----------
Then do it like this one:
----
Through: (1,2), slope 7
---------
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.
----
convert to any form you like.
==================================
Learn this.
Know it.
It will come up again and again.